# ClutterDexGrasp: A Sim-to-Real System for General Dexterous Grasping in Cluttered Scenes

### v2

@@ 1 Introduction
To enable dexterous grasping in cluttered scenes, some works first generate a non-collision grasp pose dataset specifically for cluttered scenes, then train a model to generate grasp poses and execute via motion planning [15, 14, 13]. However, the grasp poses generated by such methods are open-loop and cannot handle unexpected changes. To achieve a closed-loop grasping policy, current methods rely on real-world imitation learning (IL)[16, 17, 18] or sim-to-real reinforcement learning (RL) [19, 20, 21]. However, real-world IL requires a large amount of human-collected data to generalize across diverse cluttered scenes[16, 22, 23], which is costly and lacks scalability. Current RL-based methods, which can achieve high success rates in the real world, are mainly designed for relatively simple scenes, such as single-object tabletop grasping [11, 12]. These methods, however, aren’t directly applicable to cluttered scenes because they rely on simplified object states, like 6D poses, to ensure stable and convergent teacher policy learning, yet fail to capture the essential local geometric details crucial for dexterous grasping in cluttered scenes [24, 25]. To compactly represent the interaction between the object and the dexterous hand, [26] proposes a representation that computes the distance between the links of the dexterous hand and the graspable and ungraspable areas, which further improves teacher policy training [27].
@@
In this paper, we aim to develop a generalized closed-loop policy for dexterous target-oriented grasping in cluttered scenes, trained in a simulation environment without expert demonstrations [28, 19], and to achieve zero-shot sim-to-real transfer [29, 30, 31]. The main challenges are as follows: (1) training instability in RL due to high-dimensional observation and action spaces caused by the diversity of object geometries and layouts, as well as the high DoF of the dexterous hand [19, 21, 20], (2) the sim-to-real gap in observation space, policy safety, and physical dynamics [32, 33, 34]. To address these challenges, we propose ClutterDexGrasp, a two-stage teacher-student framework. The teacher is trained in simulation using clutter density curriculum learning to reduce the difficulty of directly learning complex grasping strategies. Additionally, we leverage the distance representation [26] and extend it to clutterd scenes by computing distance between links of the dexterous hand and the target and non-target objects, such a geometry and spatially embedded scene representation captures the local geometry of the target object and its surrounding environment, enabling the teacher policy to effectively learn human-like grasping strategies in cluttered scenes [1]. The student policy is trained using offline imitation learning with a 3D diffusion policy (DP3) [35] on partial point cloud observations, enabling zero-shot sim-to-real transfer. To ensure safety during contact-rich interactions, we introduce an interaction-aware safety curriculum that minimizes collisions and avoids harmful behaviors, such as excessive force.
@@ 2 Related Works
Transferring RL policies from simulation to real-world remains challenging for dexterous manipulation [20, 34, 33, 32, 42]. RL enables learning complex behaviors that are hard to engineer manually [28, 43, 44], but introduces sim-to-real challenges due to distributional shifts in dynamics and observations [19, 21]. Techniques like domain randomization [29, 30] improve robustness by training across randomized simulation parameters, while system identification [31, 45] calibrates simulations to better match real-world dynamics. Safety-aware RL is critical for contact-rich behaviors [46, 47, 48, 42], especially in cluttered scenes with unpredictable object interactions. Lin et al. [33] and Maddukuri et al. [32] address sim-to-real RL for dexterous manipulation, focusing on single-object tasks. Curriculum learning [38] structures training by gradually increasing task complexity while maintaining robustness and safety. Our method combines these strategies with an interaction-aware safety curriculum that progressively tightens constraints during training for safe real-world deployment.
@@
Recent advances in generative modeling have introduced diffusion models for policy learning [49, 50, 51, 52, 53, 54, 55, 42], with significant potential for modeling complex action distributions. Ze et al. [35] proposed a 3D diffusion policy that operates directly on point cloud observations, making it well-suited for robotic manipulation tasks with partial observations. Policy distillation [36] transfers knowledge from a teacher policy with privileged information to a student policy with limited observation, which is particularly valuable for sim-to-real transfer. The teacher-student framework [3, 11], combined with diffusion policies, enables zero-shot sim-to-real transfer by distilling privileged simulation knowledge into policies that operate with real-world sensory input. Our approach leverages this paradigm, using a privileged teacher policy to generate demonstrations across varying clutter densities, which are then distilled into a point-cloud-based student policy capable of zero-shot transfer to real-world environments.
@@ 3 Problem Formulation
We consider a tabletop scenario with a 7-DoF robotic arm $\mathbf{J}^{a}\in\mathbb{R}^{7}$ and a 12-DoF dexterous hand $\mathbf{J}^{h}\in\mathbb{R}^{12}$. The hand joints consist of 6-DoF actuated finger joints $\mathbf{J}^{f}\in\mathbb{R}^{6}$, and 6-DoF underactuated finger joints $\mathbf{J}^{u}\in\mathbb{R}^{6}$. The action space $\mathcal{A}\subseteq\mathbb{R}^{13}$ encompasses 7D relative changes for the arm joint $\mathbf{a}^{a}$ and 6D absolute joint positions for the actuated hand joints $\mathbf{a}^{h}$.
@@ 4.1 Teacher Policy Learning in Simulation
To enable general, dynamic, and safe grasping behavior, we train a teacher policy in simulation using a three-stage curriculum learning strategy. This curriculum progressively increases task complexity to guide policy development: (1) grasping single objects to acquire foundational manipulation skills, (2) extending to cluttered environments to learn clutter-aware behaviors, and (3) fine-tuning with safety constraints to promote robust and safe execution during real-world deployment. The policy optimization leverages a novel geometry and spatial representation with privileged state information (Sec.4.1.1). This representation encodes both object geometry and spatial relationships to enable strategic, collision-aware grasping. We detail each component below.
@@ 4.1.1 Geometry and Spatial Representation
For each finger joint, we compute distances to the nearest sampled points from both the target ($d_{pos}$) and non-target objects ($d_{neg}$), as illustrated in Fig. 9. This representation efficiently encodes both object geometry and spatial relationships between the hand and the clutter. It enables collision-aware grasping while maintaining a compact observation space that simplifies learning of complex geometric features. By incorporating this representation into both the observation space and reward function, we promote clutter-aware grasping without requiring computationally expensive point-cloud rendering. Detailed implementation is provided in Appendix B.1.
@@
$r=(c_{1}\cdot r_{\text{grasp}}+c_{2}\cdot|d_{pos}|)*(1-\bar{d}_{neg}^{\text{min}})$ (3)
where $c_{1},c_{2}>0$ are weighting coefficients, $\bar{d}_{neg}^{\text{min}}$ is the minimum absolute distance to non-target object and $r_{\text{grasp}}$ represents the base grasping reward (detailed in Appendix B.3).
4.1.2 Clutter-Density Curriculum
As the first two stages of curriculum learning, we introduce the Clutter-Density Curriculum. Given the complexity of learning general grasping in cluttered scenes, where intricate interactions and robust grasping are required, learning entirely from scratch is found to be infeasible as shown in Fig. 8. To tackle this, we decompose the learning process into two stages: (1) learning robust, generalizable grasping policy for single-object scenarios, and (2) fine-tuning the policy in cluttered environments to learn contact-rich, strategic grasping. This curriculum allows the agent to first efficiently master basic grasping skills for single objects before learning more complex behaviors required for cluttered scenes, such as moving obstacle objects away before grasping the target object. We implement both stages using a unified RL formulation with PPO [56] following the objective in Eq. (1), where policies $\pi^{E}(a|\mathcal{O}):\mathcal{O}\rightarrow\Pr(\mathcal{A})$ map privileged observations to actions. The implementation details of teacher policy learning are deferred to Appendix B.
@@ 4.1.3 Interaction Safety Curriculum
As the final stage, we introduce an interaction Safety Curriculum to ensure safe and robust grasping in cluttered scenes. This curriculum is crucial for sim-to-real transfer, as it mitigates the overly aggressive behaviors developed in the previous stages while preserving the desired human-like strategies. The safety curriculum is designed to minimize collisions and avoid excessive force during contact-rich interactions, which are common in cluttered environments.
Specifically, the policy is fine-tuned by progressively tightening a force threshold $\bar{f}$ whenever the success rate $w$ surpasses a threshold $\bar{w}$. Exceeding this force threshold incurs a sparse penalty term $r_{\text{force}}$ in the reward and leads to early termination. To further avoid colliding with table—particularly when grasping small objects, we also disabled collisions between the hand and the table in simulation, and terminate the episode if the fingertip penetrates too deeply into the table surface.
The policy is optimized with Eq. (1) with the updated reward function with safety penalties, defined as:
@@
where $c_{3}>0$ are weighting coefficients.The implementation details of safety curriculum are deferred to Appendix B.4
@@ 4.2.2 Point Cloud Alignment
To address the sim-to-real perception gap and handle occlusions in cluttered scenes, we augment the observed point cloud with synthetic, dense robot point clouds as shown in Fig.10. Similar to Dexpoint[7], our augmentation includes both arm and hand point clouds, which helps establish critical spatial relationships for accurate grasping and collision avoidance. We apply consistent point cloud preprocessing across simulation and real-world data to ensure alignment. The details are deferred to Appendix C.3.
@@ 4.2.3 System Identification for Sim-to-Real Transfer
To minimize the dynamics gap between simulation and reality, we perform system identification (SI) [31, 45] to calibrate the physical parameters of both the robotic arm and hand. The SI process iteratively compares trajectories from simulation and the real world under identical commands, optimizing simulation parameters until the simulated behavior closely matches real-world motion [34]. The details are deferred to Appendix C.4.
@@ 5.1.2 Effectiveness of Density Curriculum and Safety Curriculum
To evaluate the necessity of the Density Curriculum, we conduct an experiment by directly training the policy on cluttered scenes. Under the same training duration, training directly in clutter results in complete failure (0% success) due to complex interactions, whereas initializing from the general grasp policy achieves 87.0% success rate—demonstrating the effectiveness of progressive curriculum learning. The learning curve can be found in Fig. 8
@@ A.1 Geometric and Spatial (GS) Representation
To better understand the contribution of GS-representation, we introduce two ablation experiments with the same framework: (1) one without the proposed geometry-and-spatial representation (teacher policy w/o GS-repr) (2) and another without the distance to non-target object (teacher policy w/o negative repr).
@@ A.2 Clutter-Density Curriculum Learning
Learning dexterous grasping in cluttered scenes purely through random trial-and-error is extremely challenging. As shown in Figure 8, a policy trained from scratch under identical settings fails to make progress, with success rates remaining at zero. To address this, we introduce a two-stage Clutter-Density Curriculum: the policy is first trained on general single-object grasping, then fine-tuned in cluttered scenes to develop strategic, human-like behaviors. This staged approach enables effective learning, as illustrated by the performance curve in Figure 8.
Figure 8: Learning curves of the cluttered-scene policies (1) Teacher Policy (w/o safety): initialized with stage 1 general single-object grasping policy, (2) w/o curriculum: trained from scratch directly in cluttered scenes.
@@ B.1 Geometric and Spatial (GS) Representation
Direct grasp attempts in dense cluttered scenes often result in unsafe collisions with surrounding non-target objects, we therefore introduced a Geometric and Spatial Representation to encourage collision-minimized behavior. At each time step, $N_{\text{pos}}=200$ points are sampled from the target object and $N_{\text{neg}}=50$ points from surrounding objects using their mesh files. For each finger joint, distances to the nearest sampled points from both target ($d_{pos}$) and non-target objects ($d_{neg}$) are computed to form the representation, as shown in Fig. 9. This representation is embedded into observation space and reward function, providing essential spatial and geometric information for clutter-aware grasping, while bypassing the need for time-consuming point-cloud rendering.
@@
Figure 9: Visualization of the Geometric and Spatial Representation. For each finger joint, distances to the nearest $N_{\text{pos}}$ surface points sampled from the target object mesh ($d_{\text{pos}}$, green) and $N_{\text{neg}}$ points from surrounding non-target object meshes ($d_{\text{neg}}$, red) are computed and visualized.
@@ B.3 Teacher Reward Function
At each timestep, the complete reward function is defined as:
$r=\left(c_{1}\cdot r_{grasp}+c_{2}\cdot|d_{pos}|\right)\cdot(1-\bar{d}_{neg}^{\text{min}})$ (5)
Here, $\bar{d}_{neg}^{\text{min}}$ denotes the minimum absolute distance to any non-target object (negative object) observed over the entire grasping episode. It serves as a global penalty term to discourage risky proximity to non-target objects at any point during execution, and $r_{grasp}$ is defined as:
@@
$+c_{6}\cdot|d_{mid}|\big{)}$ middle point reward
where $c_{1},c_{2},c_{4},c_{5},c_{6}>0$ are weighting coefficients. $d_{mid}$ is the distance between middle-point between index finger and thumb and center of object. $p_{current}$ and $p_{goal}$ denote the current position of the target object and the desired position the target object should be lifted to after a successful grasp.
@@
$r_{stage1}=c_{1}\cdot r_{grasp}+c_{2}\cdot|d_{pos}|$ (7)

### v3

@@ 1 Introduction
To enable dexterous grasping in cluttered scenes, some works first generate a non-collision grasp pose dataset specifically for cluttered scenes, then train a model to generate grasp poses and execute via motion planning [15, 14, 13]. However, such methods are open-loop and cannot handle unexpected changes. To achieve a closed-loop grasping policy, current methods rely on real-world imitation learning (IL)[16, 17, 18] or sim-to-real reinforcement learning (RL) [19, 20, 21]. However, real-world IL requires a large amount of human-collected data to generalize across diverse cluttered scenes[16, 22, 23], which is costly and lacks scalability. Current RL-based methods, which can achieve high success rates in the real world, are mainly designed for relatively simple scenes, such as single-object tabletop grasping [11, 12]. These methods, however, aren’t directly applicable to cluttered scenes because they rely on simplified object states, like 6D poses, to ensure stable and convergent teacher policy learning, yet fail to capture the essential local geometric details crucial for dexterous grasping in cluttered scenes [24, 25]. To compactly represent the interaction between the object and the dexterous hand, [26] proposes a representation that computes the distance between the links of the dexterous hand and the graspable and ungraspable areas, which further improves teacher policy training [27].
@@
In this paper, we aim to develop a generalized closed-loop policy for dexterous target-oriented grasping in cluttered scenes, trained in a simulation environment without expert demonstrations [28, 19], and to achieve zero-shot sim-to-real transfer [29, 30, 31]. The main challenges are as follows: (1) training instability in RL due to high-dimensional observation and action spaces caused by the diversity of object geometries and layouts, as well as the high DoF of the dexterous hand [19, 21, 20], (2) the sim-to-real gap in observation space, policy safety, and physical dynamics [32, 33, 34]. To address these challenges, we propose ClutterDexGrasp, a two-stage teacher-student framework. The teacher is trained in simulation using clutter density curriculum learning to reduce the difficulty of directly learning complex grasping strategies. Additionally, we leverage the distance representation [26] and extend it to clutterd scenes by computing distance between links of the dexterous hand and the target and non-target objects, such a geometry and spatially embedded scene representation captures the local geometry of the target object and its surrounding environment, enabling the teacher policy to effectively learn human-like grasping strategies in cluttered scenes [1]. To ensure safety during contact-rich interactions, we introduce an interaction-aware safety curriculum that minimizes collisions and avoids harmful behaviors, such as excessive force. The student policy is trained using offline imitation learning with a 3D diffusion policy (DP3) [35] on partial point cloud observations, enabling zero-shot sim-to-real transfer.
@@ 2 Related Works
Transferring RL policies from simulation to real-world remains challenging for dexterous manipulation [20, 34, 33, 32, 42]. RL enables learning complex behaviors that are hard to engineer manually [28, 43, 44], but introduces sim-to-real challenges due to distributional shifts in dynamics and observations [19, 21]. Techniques like domain randomization [29, 30] improve robustness by training across randomized simulation parameters, while system identification [31, 45] calibrates simulations to better match real-world dynamics. Safety-aware RL is critical for contact-rich behaviors [46, 47, 48, 42], especially in cluttered scenes with unpredictable object interactions. [33] and [32] address sim-to-real RL for dexterous manipulation, focusing on single-object tasks. Curriculum learning [38] structures training by gradually increasing task complexity while maintaining robustness and safety. Our method combines these strategies with an interaction-aware safety curriculum that progressively tightens constraints during training for safe real-world deployment.
@@
Recent advances in generative modeling have introduced diffusion models for policy learning [49, 50, 51, 52, 53, 54, 55, 42], with significant potential for modeling complex action distributions. [35] proposed a 3D diffusion policy that operates directly on point cloud observations, making it well-suited for robotic manipulation tasks with partial observations. Policy distillation [36] transfers knowledge from a teacher policy with privileged information to a student policy with limited observation, which is particularly valuable for sim-to-real transfer. The teacher-student framework [3, 11], combined with diffusion policies, enables zero-shot sim-to-real transfer by distilling privileged simulation knowledge into policies that operate with real-world sensory input. Our approach leverages this paradigm, using a privileged teacher policy to generate demonstrations across varying clutter densities, which are then distilled into a point-cloud-based student policy capable of zero-shot transfer to real-world environments.
@@ 3 Problem Formulation
We consider a tabletop scenario with a 7-DoF robotic arm $\mathbf{J}^{a}\in\mathbb{R}^{7}$ and a 12-DoF dexterous hand $\mathbf{J}^{h}\in\mathbb{R}^{12}$. The hand joints consist of 6-DoF actuated finger joints $\mathbf{J}^{f}\in\mathbb{R}^{6}$, and 6-DoF underactuated finger joints $\mathbf{J}^{u}\in\mathbb{R}^{6}$. The action space $\mathcal{A}\subseteq\mathbb{R}^{13}$ encompasses 7D relative changes for the arm joint $\mathbf{a}^{a}$ and 6D absolute joint positions for the actuated hand joints $\mathbf{a}^{h}$ as [56].
@@ 4.1 Teacher Policy Learning in Simulation
To enable general, dynamic, and safe grasping behavior, we train a teacher policy in simulation using a three-stage curriculum learning strategy. The first two curriculum progressively increases task complexity to guide policy development. The final stage fine-tunes the policy with safety constraints to promote robust and safe execution during real-world deployment. We implement three stages using a unified RL formulation with PPO [57] and a novel geometry and spatial representation with privileged state information (Sec.4.1.1). Following the objective in Eq. (1), policies $\pi^{E}(a|\mathcal{O}):\mathcal{O}\rightarrow\Pr(\mathcal{A})$ map privileged observations to actions. We detail each component below.
@@ 4.1.1 Geometry and Spatial Representation
For each finger link, we compute the 3D distance vector to the nearest sampled points from both the target ($d_{pos}$) and non-target objects ($d_{neg}$), as illustrated in Fig. 9. This representation efficiently encodes both object geometry and spatial relationships between the hand and the clutter. It enables collision-aware grasping while maintaining a compact observation space that simplifies learning of complex geometric features and avoiding computationally expensive point-cloud rendering. Detailed implementation is provided in Appendix B.1.
@@
$r=(c_{1}\cdot r_{\text{grasp}}+c_{2}\cdot r_{pos})\cdot(1-r_{neg})$ (3)
where $c_{1},c_{2}>0$ are weighting coefficients. Here, $r_{pos}$ provides a reward term encouraging proximity to the target, while $r_{neg}$ imposes a penalty for risky closeness to non-target objects, and $r_{\text{grasp}}$ represents the base grasping reward (detailed in Appendix B.3).
4.1.2 Two-Stage Clutter-Density Curriculum
As the first two stages of curriculum learning, we introduce the Clutter-Density Curriculum. Given the complexity of learning general grasping in cluttered scenes, where intricate interactions and robust grasping are required, learning entirely from scratch is found to be infeasible as shown in Fig. 8. To tackle this, we decompose the learning process into two stages: (1) learning robust, generalizable grasping policy for single-object scenarios, and (2) fine-tuning the policy in cluttered environments to learn contact-rich, strategic grasping. This curriculum allows the agent to first efficiently master basic grasping skills for single objects before learning more complex behaviors required for cluttered scenes, such as moving obstacle objects away before grasping the target object.
@@ 4.1.3 Interaction Safety Curriculum
As the final stage, we introduce an interaction Safety Curriculum to ensure safe and robust grasping in cluttered scenes. This curriculum is crucial for sim-to-real transfer, tempering overly aggressive behaviors while preserving human-like strategies. It is designed to minimize collisions and excessive force during contact-rich interactions common in cluttered environments. Specifically, the policy is fine-tuned by progressively tightening a force threshold $\bar{f}$ whenever the success rate $w$ surpasses a threshold $\bar{w}$. Exceeding this force threshold incurs a sparse penalty term $r_{\text{force}}$ in the reward and leads to early termination. The reward function Eq. (7) is updated with safety penalties as:
@@
where $c_{3}>0$ are weighting coefficients. To avoid colliding with table, particularly when grasping small objects, we also disabled collisions between the hand and the table in simulation, and terminate the episode if the fingertip penetrates too deeply into the table surface. The implementation details of safety curriculum are deferred to Appendix B.4.
@@ 4.2.2 Point Cloud Alignment
To address the sim-to-real perception gap and handle occlusions in cluttered scenes, we augment the observed point cloud with synthetic, dense robot point clouds as shown in Fig.10. Similar to Dexpoint[7], our augmentation includes both arm and hand point clouds, which helps establish critical spatial relationships for accurate grasping and collision avoidance. The details are deferred to Appendix C.3.
@@ 4.2.3 System Identification for Sim-to-Real Transfer
To minimize the dynamics gap between simulation and reality, we perform system identification (SI) [31, 60] to calibrate the physical parameters of both the robotic arm and hand. The SI process iteratively compares trajectories from simulation and the real world under identical commands, optimizing parameters until behaviors closely match  [34], detailed in Appendix C.4.
@@ 5.1.2 Effectiveness of Density Curriculum and Safety Curriculum
To evaluate the necessity of the Density Curriculum, we conduct an experiment by directly training the policy on cluttered scenes. Under the same training duration, training directly in clutter results in complete failure (0% success) due to complex interactions, whereas initializing from the general grasp policy achieves 87.0% success rate—demonstrating the effectiveness of progressive curriculum learning. The learning curve can be found in Fig. 8.
@@ 7 Limitation
Acknowledgments
We thank Tianyu Wang, Hongwei Fan and Hang Dai for their support in conducting real-world experiments, Hongjie Fang and Tyler Ga Wei Lum for their insightful discussion. We are also grateful to Yanzhou Jin and Haotian Jin for their assistance with the hardware in this project. This research was supported by The National Youth Talent Support Program (8200800081) and National Natural Science Foundation of China (62376006).
@@ A.1 Geometric and Spatial (GS) Representation
To better understand the contribution of GS-representation, we introduce two ablation experiments with the same framework: (1) one excludes the proposed geometry-and-spatial representation from both observation and reward function (teacher policy w/o GS-repr) (2) and another without the distance to non-target object (teacher policy w/o negative repr).
@@ A.2 Clutter-Density Curriculum Learning
Learning dexterous grasping in cluttered scenes purely through random trial-and-error is extremely challenging. As shown in Figure 8, a policy trained only on clutter scenes with identical settings and full two-stage duration (w/o curriculum) fails to make progress, with success rates remaining at zero. To address this, we introduce a two-stage Clutter-Density Curriculum: the policy is first trained on general single-object grasping, then fine-tuned in cluttered scenes to develop strategic, human-like behaviors. This staged approach enables effective learning, as illustrated by the performance curve in Figure 8.
Figure 8: Learning curves of the cluttered-scene policies (1) Teacher Policy (w/o safety): initialized with stage 1 general single-object grasping policy, (2) w/o curriculum: trained from scratch directly in cluttered scenes for the full two-stage duration.
@@ B.1 Geometric and Spatial (GS) Representation
Direct grasp attempts in dense cluttered scenes often result in unsafe collisions with surrounding non-target objects, we therefore introduced a Geometric and Spatial Representation to encourage collision-minimized behavior. This representation serves as a privileged observation for the teacher policy, providing essential spatial and geometric information for clutter-aware grasping, while bypassing the need for time-consuming point-cloud rendering. At each time step, we have $N_{\text{env}}$ environments. For each environment, we sample 200 points from a randomly selected target object, and represent all points as a $N_{\text{env}}\times 200\times 3$ matrix. We sample 50 points from non-target objects $N_{\text{neg}}$, and represent all points as a $N_{\text{env}}\times 50\times 3$ matrix. We get 11 base positions from selected finger links, and represent as a $N_{\text{env}}\times 11\times 3$ matrix. Using torch.norm, we compute distances between link positions and both target object points and non-target object points, identify nearest points, and construct 3D vectors from link bases: $d_{pos},d_{neg}\in\mathbb{R}^{N_{\text{env}}\times 11\times 3}$, which serve as observations. During training, $N_{\text{env}}=16384$ with computation time cost 6ms. This computation is not required by the student policy during real-world deployment.
@@
Figure 9: Visualization of the Geometric and Spatial Representation. For each finger joint, distances to the nearest $N_{\text{pos}}$ surface points sampled from the target object mesh ($d_{pos}$, green) and $N_{\text{neg}}$ points from surrounding non-target object meshes ($d_{neg}$, red) are computed and visualized.
@@ B.3 Teacher Reward Function
We first introduce the definitions of $r_{pos}$ and $r_{neg}$ as:
$r_{pos}=\exp{(-\alpha_{pos}\cdot\lVert d_{pos}\rVert_{2})}$ (5)
$r_{neg}=\exp{(-\alpha_{neg}\cdot\bar{d}_{neg}^{\text{min}})}$ (6)
Here, $r_{pos}$ provides a positive reward that encourages the dexterous hand to approach the target object. Conversely, $r_{neg}$ introduces a global penalty term based on $\bar{d}_{neg}^{\text{min}}$, the minimum absolute distance to any non-target (negative) object surface observed throughout the entire grasping episode. This term serves as a global penalty term to discourage risky proximity to non-target objects at any point during execution. At each timestep, the complete reward function is defined as:
$r=\left(c_{1}\cdot r_{grasp}+c_{2}\cdot r_{pos}\right)\cdot(1-r_{neg})$ (7)
and $r_{grasp}$ is defined as:
@@
$+c_{6}\cdot\exp{(-\alpha_{mid}\cdot\lVert d_{mid}\rVert_{2})}\big{)}$ middle point reward
where $c_{1},c_{2},c_{4},c_{5},c_{6},\alpha_{pos},\alpha_{neg},\alpha_{mid}>0$ are weighting coefficients. $d_{mid}$ is the 3D vector from middle-point between index finger and thumb to center of object. $p_{current}$ and $p_{goal}$ denote the current position of the target object and the desired position the target object should be lifted to after a successful grasp.
@@
$r_{stage1}=c_{1}\cdot r_{grasp}+c_{2}\cdot r_{pos}$ (9)